# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

### v2

@@ Abstract
Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN). While recent advancements have yielded promising outcomes, they primarily rely on RGB images for environmental representation, often overlooking the underlying semantic knowledge and spatial cues. Intuitively, humans inherently ground textual semantics within the spatial layout during indoor navigation. Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to facilitate navigation. SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in the agent’s immediate surroundings. Additionally, a Depth-based Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts. Experimental results demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON). The source code are publicly available at https://github.com/HCI-LMC/VLN-SUSA.
@@ 1 Introduction
With the advancement of multimodal technology [43, 32, 63, 47, 24], embodied VLN tasks [3, 40, 35, 37, 7, 44] have garnered significant attention due to their promising applications in robotics and intelligent assistance. Among them, conventional VLN tasks [3, 22, 20] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [65, 40, 45] demand agents to identify predefined objects based on high-level instructions. Consequently, agents must profoundly cognize and exploit the rich environmental information and ground textual instructions to brilliantly accomplish these arduous navigation tasks.
@@
A significant body of VLN research has focused on Transformer-based models [4, 18, 38, 55, 39] for interacting with instructions, environmental images, and historical trajectories. While these models have undergone substantial development and exhibit efficacy, two key challenges remain in VLN tasks. First, VLN agents struggle to precisely match the landmarks (e.g., “fireplace” in Fig. 1) in instructions with the entities in environments, due to the intrinsic modality heterogeneity between vision and language. To address this, recent work [11, 54, 28] has leveraged contrastive learning to promote alignment, while others have developed repositories of candidate object concepts [53, 29] or utilized large language models (LLMs) to identify the most relevant landmarks in the visual scene [30]. However, these methods often lack an explicit understanding of environments from the perspective of textual semantics (i.e., the linguistic descriptions of landmarks). Intuitively, when navigating indoors or searching for objects, humans naturally associate the textual semantics of their surroundings.
The second challenge pertains to the scarcity of diverse environmental sources, which often leads VLN agents to overfit to familiar environments, hindering generalization. To alleviate this limitation, various methods [23, 57, 17, 6, 50, 27] have employed data augmentation to increase the environmental quantity and diversity. More recent endeavors enhance spatial perception by constructing finer-grained graph [25, 2], semantic map [15, 19], grid [58, 33], or volumetric [34] representations. However, these methods predominantly rely on RGB images, overlooking the depth images, which provides more intuitive spatial structure, especially in stereoscopic navigation environments. Furthermore, few approaches [48] treat depth as a distinct modality capable of offering complementary spatial information, instead using it merely as an auxiliary signal to refine RGB-based representations. In conclusion, the co-occurrence semantic and spatial environment representations, beyond surface RGB images, has not been fully explored or exploited.
@@
In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, designed to explicitly capture both semantic and spatial information during navigation. As illustrated in Fig. 1, we first construct a textual semantic panorama by extracting salient landmarks as the textual semantics for each view to explicitly match the textual semantics with language instructions, and select the most instruction-relevant view for navigation prediction, thereby bridging the semantic gap between the visual and linguistic modalities. To enhance the spatial awareness, we then progressively construct an independent depth exploration map, which helps mitigate the risk of the agent overfitting to the information-dense RGB images. To enabling the agent to absorb complementary information from different environmental perspectives, we adaptively fuse these hybrid semantic-spatial environmental representations to facilitate informed navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions. We empirically substantiate that exclude semantic panoramas and depth exploration maps during pretraining manifest markedly enhanced generalization capabilities. Experimental results demonstrate that our agent achieves significantly improved navigation performance, outperforming state-of-the-art methods across three vision-and-language navigation benchmarks: R2R [3], REVERIE [40], and SOON [65].
@@
  • SUSA devises a textual-aware semantic understanding module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
  • SUSA presents a depth-based spatial perception module to explore and represent environmental layouts beyond surface RGB content.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). Vision-and-Language Navigation (VLN) tasks [3, 40, 65] require an agent to navigate sequentially towards a designated location or object based on natural language instructions. Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [56, 52, 66, 13, 21, 36] primarily relied on reinforcement and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [64, 56] designs. More recently, to better generalize to unknown environments and ground instructions, many researchers [4, 18, 38, 55, 39] have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories. While these approaches have propelled VLN forward, the co-occurrence environmental sources have not been fully utilized. This paper introduces text semantics and depth information to enrich environmental representations.
@@
Semantic Alignment in VLN. To align the relationship between the environment and instructions, current researchers and practitioners in the VLN community can be divided into two main categories: implicit and explicit approaches. Implicit methods [11, 54, 28] commonly employ contrastive learning to associate visual contexts with textual instructions or use attention mechanisms to model the relationship. However, due to the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities. Explicit alignment techniques aim to project both semantics into a unified modality space. For example, some methods transform the environment into a text-based modality, Console [30] scores landmark co-occurrences to identify key navigation objects, while AACL [29] leverages CLIP [43] to refine textual entities, compacting them with visual features. However, these explicit methods often directly “stitch” textual and visual environment representations, which may leading to an overreliance on precise but abstract text while underutilizing rich yet redundant visual data. We thus use environmental textual semantic as an independent modality to explicitly match the most instruction-relevant landmarks.
Environment Representation in VLN. Limited training data and embodied environment scarcity in VLN tasks usually hinder the agent to generalize the unknown scenes or layouts. Previous approaches mitigate this issue through data augmentation [31, 57] or by enhancing visual representations[61, 34]. ScaleVLN [57] introduces thousands of photorealistic environments and FDA [17] incorporates low-frequency interference images to emphasize essential high-frequency details. Advanced visual representations, such as topological maps and grid-based layouts, further enable agents to retain temporal navigation paths and more precisely perceive spatial structures [1, 58]. For example, DUET [5] employs dual-scale transformers [51] to balance global and local decision-making, while VER [34] adopts volumetric models to represent 3D spatial structures. While most research relies on RGB representations, which tend to capture redundant appearance features, our method incorporates depth data, offering agents a more intuitive and spatially aware understanding of the environment.
@@
Figure 2: The detailed architecture of our proposed SUSA model. The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively. The Hybrid Representation Fusion (HRF) module is designed for aligning environmental representations with instructions.
@@ 3.1 Preliminary
Task Formulation. We focus on discrete indoor VLN tasks [3, 40, 65] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths. At the outset of each episode, the agent, equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words. We use BERT [9] to extract instruction features and then embed them as =$\mathcal{I}=\{w_{1},\dotso,w_{l}\}$. At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views. During navigation, the agent must decide its next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$. The sequential decision-making process is iterated until the agent decides to stop at the current location or exceeds the maximum navigation steps. For goal-oriented tasks [40, 65], the agent is also tasked with locating a target object.
Overview of the Proposed Method. Our objective is to equip the agent with a textual semantic panorama and a depth exploration map to accord complementary environmental content. We adopt DUET [5] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions. The SUSA framework consists of three main modules: (a) textual-aware semantic selection is detailed in Sec. 3.2, (b) depth-aware visual environment exploration is presented in Sec. 3.3, and (c) hybrid representation fusion and action prediction is subsequently introduced in Sec. 3.4, as illustrated in Fig. 2.
@@ 3.2 Textual-Aware Semantic Understanding
Salient Landmarks Extractor. The pretrained visual-language model readily provides the capability for agents to perceive textual semantics by generating detailed captions from RGB images. As illustrated in Fig. 2, we utilize off-the-shelf BLIP-2-FlanT5-xxL [24], a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”. Following [14, 23], this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks. These descriptions typically involve room types and objects (e.g.,“a living room with two chairs”), which assist in the object recognition. We further discuss the quality of the generated descriptions in supplementary materials. Additionally, we employ the CLIP (ViT-L-14-336px) text encoder [43] to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$. During navigation, we leverage the frozen BLIP-2 and CLIP models to empower the agent to obtain real-time textual semantic features from the navigable nodes in its vicinity.
Textual Semantic Selector. Previous methods [53, 29] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed. In contrast, we treat the textual semantic as a distinct modality, directly matching salient landmarks with the instructions. As shown in Fig. 2 and Fig. 3, our textual semantic selector combines both $static$ and $dynamic$ matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
@@
$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{S}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
@@
where the balance factor $\delta$ regulates the relative contributions of each matching strategy. By mapping the environment to the same textual modality, the textual semantic selector enables the agent to better understand the environment context from a textual semantic perspective, thus facilitating grounded language instructions.
3.3 Depth-Aware Spatial Perception
@@
Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited. To enhance the spatial perception for agent, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [5].
@@
Depth-Based Exploration Mapping. Concretely, we obtain ground truth depth images for each view from the Matterport3D simulator. As illustrated in Fig. 2, we then employ CLIP (ViT-L/14@336px) and ResNet-50 [16] (pretrained on the Gibson [59]) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view. Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images. The panorama encoding modules for both modalities share weights, since each view in the depth and RGB panoramas is accordant. Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel. The explored depth and RGB trajectory maps are formalized as $\mathcal{T}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node. These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable points when considering potential backtracking.
@@
Cross-Modal Interaction and Reasoning. To defer to instructions, we utilize a multi-layer cross-encoder that interacts with both the language instructions and the depth exploration map features to generate the refined depth map $\tilde{\mathcal{T}^{\mathcal{D}}}$. This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention [5] (GASA), cross-modal attention, and a feed-forward network. The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation points, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices. In parallel, we facilitate similar cross-modal interactions between the language instructions and the RGB exploration map $\mathcal{T}^{\mathcal{R}}$, as well as the RGB panoramas $\mathcal{X}^{\mathcal{R}}$. These are encoded as $\tilde{\mathcal{T}^{\mathcal{R}}}$ and $\tilde{\mathcal{X}^{\mathcal{R}}}$, thereby enabling the model to perform both global and local predictions based on the given instructions $\mathcal{I}$. Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises and biases.
@@
Figure 4: Illustration of the proposed HRF module in  Fig. 2.
3.4 Hybrid Representation Fusion and Prediction

### v4

@@ Abstract
Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN). Existing approaches primarily rely on RGB images for environmental representation, underutilizing latent textual semantic and spatial cues and leaving the modality gap between instructions and scarce environmental representations unresolved. Intuitively, humans inherently ground semantic knowledge within spatial layouts during indoor navigation. Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to encourage agents to ground environment from diverse perspectives. SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in agent’s immediate surroundings. Additionally, a Depth-enhanced Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts. Experiments demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON). The source code will be publicly available.
@@ 1 Introduction
With the advancement of multimodal technology [44, 47, 29, 59], several embodied VLN tasks [2, 41, 37, 8] have garnered significant attention due to their promising applications in robotics and intelligent assistance. Among them, conventional VLN tasks [2, 27] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [64, 41, 45] demand agents to identify predefined objects based on high-level instructions. Consequently, VLN agents are required to profoundly cognize the spatial information of environment and ground them with textual instructions to brilliantly accomplish arduous navigation tasks.
@@
Considerable literature has exhibited the efficacy of Transformer-based models [5, 20, 38, 52, 40] for interacting with instructions, RGB-based environmental observations, and historical trajectories. However, we suggest that RGB-only visual environment representations leave VLN fraught with two challenges: the modality gap between instructions and the environment, and the scarcity of training environmental representations. Firstly, VLN agents struggle to precisely match the landmarks in language instructions with the entities in visual environments, due to the intrinsic modality heterogeneity. To address this, recent works [12, 32] utilized contrastive learning to promote alignment, while others leveraged candidate object concepts [50, 33], large language models (LLMs) [34] or retrieved knowledge [31] to explicitly reason the textual semantics (i.e., linguistic descriptions about landmarks) to enhance environmental representations. However, these methods do not directly align textual semantics with landmarks (e.g., “fireplace” in Fig. 1) in instruction, yielding trivial gain.
The second challenge pertains to the monotonous and limited environmental representations, which often leads agents to overfit to familiar spatial layouts [51]. To alleviate this limitation, recent endeavors [55, 7] augment the quantity of environmental scenes, while others diversify environmental representations by constructing graphs [16, 23], grids [56, 1], or volumetric environments [36]. However, these RGB-based methods easily suffer from attention imbalance or environmental bias [65, 61]. FDA [18] utilizes frequency information to enrich environmental representations. However, depth images vividly convey spatial information (e.g., doors and corridors), which is often overlooked in discrete navigation tasks [49]. Few works [1, 56] incorporate depth, they merely project depth values onto the corresponding spatial locations of RGB images, rather than independently grounding depth[53]. Overall, the underlying textual semantic and spatial representations remains underutilized, though studies [22] suggest that injecting inductive biases where different modules access distinct modalities, offers a promising solution.
@@
In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, which individually grounds semantic and spatial representations of the environment with instructions and synergistically predicts navigation actions from diverse perspectives. As illustrated in Fig. 1, we first construct a view-level textual panorama by extracting salient landmarks as the textual semantics to explicitly match them with language instructions, and select the most instruction-relevant view, thereby bridging the semantic gap between the visual and linguistic modalities. Further, beyond RGB images, we progressively construct an independent trajectory-level depth exploration map to enhance the agent’s spatial perception of the environment, also directly aligning it with instructions. To enable the agent to absorb complementary information from different environmental perspectives, we aggregate these hybrid semantic-spatial environmental representations to synergistically facilitate hierarchical navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions. We empirically substantiate that exclude semantic panoramas and depth exploration maps during pretraining manifest markedly enhanced generalization capabilities. Experimental results demonstrate that our agent achieves significantly improved navigation performance, outperforming state-of-the-art methods across three vision-and-language navigation benchmarks: R2R [2], REVERIE [41], and SOON [64].
@@
  • SUSA devises a textual-aware semantic understanding (TSU) module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
  • SUSA further presents a depth-enhanced spatial perception (DSP) module to explore and represent environmental layouts beyond surface RGB content.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). Vision-and-Language Navigation (VLN) tasks [2, 41, 64] require an agent to navigate in embodied environments sequentially towards a designated location or object based on natural language instructions. Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [54, 66, 14, 26] primarily relied on reinforcement learning and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [63, 54] designs. More recently, to better generalize to unknown environments and ground instructions, many researchers [5, 20, 38, 52, 40] have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories. While these approaches have propelled VLN forward, they mainly rely on fragile RGB-based environment representations. This paper encourages the agent to perceive and ground the environment from text semantics and depth perspectives.
@@
Semantic Alignment in VLN. To align the semantic relationship between the environment and instructions, current practitioners in the VLN community can be divided into two categories: implicit and explicit approaches. Implicit methods [12, 51, 32] commonly employ contrastive learning or attention mechanisms to associate visual contexts with textual instructions. However, given the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities. Explicit alignment techniques aim to project both semantics into a same modality space. For example, some methods transform the environment into a text-based modality, Console [34] scores landmark co-occurrences to identify key objects, and AACL [33] leverages CLIP [44] to refine textual entities. Similar to our approach, KERM [31] utilizes retrieved knowledge as textual semantics for each view. However, these explicit methods directly “stitch” textual and visual environment representations, which may lead to an overreliance on text modality yet underutilizing visual data [67]. We thus take environmental textual semantic as an independent modality to directly match the most instruction-relevant landmarks.
Environment Representations in VLN. Limited training sources and environmental representations in VLN tasks hinder the agent to generalize the unknown layouts. Prior endeavors mitigate this limitation through data augmentation [35, 55] or by diversifying visual representations [60, 1, 56, 36]. However, these approaches primarily rely on RGB images, while alternative modalities such as frequency [18], spectral[24], depth [49] and semantic information [53, 21, 4] remain relatively underexplored. Among them, SEAT [53] enhances cross-domain environmental perception by querying the Matterport3D simulator[3] for depth and semantics. However, SEAT fuses these multi-type representations into a unified space without explicitly grounding each modality. Moreover, these semantic representations [53, 21] are not inherently textual, limiting their ability to bridge the modality gap. In contrast, our method constructs the view-level textual panoramas and trajectory-level depth exploration maps, directly grounding them with instructions to perceive rich environmental representations.
@@
Figure 2: The detailed architecture of our proposed SUSA model. The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively. The Hybrid Representation Aggregation (HRA) module is designed for aligning environmental representations with instructions.
@@ 3.1 Preliminary
Task Formulation. We focus on discrete VLN tasks [2, 41, 64] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths. At the outset of each episode, the agent equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words. We use BERT [10] to extract instruction features and then embed them as =$\mathcal{I}=\{w_{1},\dotso,w_{l}\}$. At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views. During navigation, the agent make the next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$. The sequential decision-making process is iterated until the agent decides to stop at the current location or exceeds the maximum navigation steps. For goal-oriented tasks [41, 64], the agent is also tasked with locating a target object.
Overview of the Proposed Method. Our objective is to accord complementary environmental content by equipping the agent with a textual semantic panorama and a depth exploration map. Following prevailing approaches [18, 30, 19], we adopt DUET [6] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions. The SUSA framework consists of three main modules: (a) textual-aware semantic selection is detailed in Sec. 3.2, (b) depth-aware visual environment exploration is presented in Sec. 3.3, and (c) hybrid representation aggregation and hierarchical action prediction is subsequently introduced in Sec. 3.4, as illustrated in Fig. 2.
@@ 3.2 Textual-Aware Semantic Understanding
Salient Landmarks Extractor. The pretrained visual-language model readily provides the capability for agents to perceive textual semantics by generating detailed captions from RGB images. As illustrated in Fig. 2, we utilize off-the-shelf BLIP-2-FlanT5-xxL [29], a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”. Following [15, 28], this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks. These descriptions typically involve room types and objects (e.g.,“a living room with two chairs”), which assist in the object recognition. We further discuss the quality of the generated descriptions in supplementary materials. Additionally, we employ the CLIP (ViT-L-14-336px) text encoder [44] to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$. Thanks to the powerful yet frozen BLIP-2 and CLIP models, we empower the agent to obtain textual semantic features from the navigable nodes in its vicinity during navigation.
Textual Semantic Selector. Previous methods [50, 33, 31] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed. Similar to [39], we treat the textual semantic as a perceptual representation for navigation in this module, directly matching salient landmarks with the instructions. As shown in Fig. 2 and Fig. 3, our textual semantic selector combines both $static$ and $dynamic$ matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
@@
$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{X}^{\mathcal{S}}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
@@
where the balance factor $\delta$ regulates the relative contributions of each matching strategy. By mapping the environment to the same textual modality, the textual semantic selector enables the agent to explicitly understand the environment context from a textual semantic perspective, thus facilitating grounded language instructions.
3.3 Depth-Enhanced Spatial Perception
@@
Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited in earlier works. To enhance the spatial perception, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [6].
@@
Depth-Enhanced Exploration Mapping. Concretely, we obtain ground truth depth images for each view from the Matterport3D simulator. As illustrated in Fig. 2, we then employ ResNet-50 [17] (pretrained on the Gibson [57]) and CLIP (ViT-L/14@336px) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view. Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images. The panorama encoding modules for both modalities share weights, since each view in the depth and RGB panoramas is accordant. Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel. The explored depth and RGB trajectory maps are formalized as $\mathcal{T}_{t}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node. These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable nodes when considering backtracking.
@@
Cross-Modal Interaction and Reasoning. To defer to instructions, we utilize a multi-layer cross-encoder that interacts with both the language instructions and the depth exploration map features to generate the refined depth map $\tilde{\mathcal{T}^{\mathcal{D}}}$. This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention (GASA) [6], cross-modal attention, and a feed-forward network. The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation nodes, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices. In parallel, we facilitate similar cross-modal interactions between the language instructions and the RGB exploration map $\mathcal{T}^{\mathcal{R}}$, as well as the RGB panoramas $\mathcal{X}^{\mathcal{R}}$. These are encoded as $\tilde{\mathcal{T}^{\mathcal{R}}}$ and $\tilde{\mathcal{X}^{\mathcal{R}}}$, thereby enabling the model to perform both global and local predictions based on the given instructions $\mathcal{I}$. Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises [25] or biases [22, 51].
@@
Figure 4: Illustration of the pipeline for fusing hybrid environmental representations in the proposed HRA module in Fig. 2.
3.4 Hybrid Representation Aggregation